An aeroplane of mass 17000kg lands with a velocity of 75m/s reverse thrust from the engines slows it to a speed of 25m/s in a distance of 230m. What is the size of the reverse thrust supplied by the engines
1
Expert's answer
2019-11-11T15:52:48-0500
"V^2-v^2=2as"

"\left(\frac{75}{3.6}\right)^2-\left(\frac{25}{3.6}\right)^2=2(230)a"

"a=0.8387\frac{m}{s^2}"

"F=ma=(17000)(0.8387)=14258\ N"